OpenStudy (dboyette):

once again I am not getting this. assume that the population of heights of male college students is normally distibuted with a mean of 69.09 and standard deviation of 4.71. A random sample of 92 heights is obtained. find the mean and standard error of the x distribution. find P(x>68.5)

OpenStudy (amistre64):

standard error is another term for standard deviation

OpenStudy (amistre64):

z = (68.5-69.09)(sqrt(92))/4.71 = -1.2

OpenStudy (dboyette):

isn;t the error 1/2 of the mean?

OpenStudy (dboyette):

i mean the deviation?

OpenStudy (amistre64):



OpenStudy (amistre64):

no the deviation is given in the problem

OpenStudy (dboyette):

so we would do this like the last one, right? z=x-o/n/sqrt

OpenStudy (amistre64):

we can yes, but a denom IN the denom is just a multiplier

OpenStudy (amistre64):

z=x-u/o~ * sqrt(n)

OpenStudy (amistre64):

for a ztable we can use this http://www.pavementinteractive.org/index.php?title=Normal_Distribution

OpenStudy (amistre64):

at -1.2 we have an area of .1151 but we want the area to the right of the z, not the left. 1 - .1151 = .8839

OpenStudy (amistre64):

well, .8849

OpenStudy (amistre64):